Rosenthal Company manufactures bowling balls through two processes: Molding and Packaging. In the Molding Department, the ureYour answer is correct. Prepare a schedule showing physical units of production Physical units Units to be accounted for Work✓ Your answer is correct. Compute the unit costs of production. Materials Conversion Costs Total Unit Cost Unit Costs $ 17 AtPrepare a production cost report for the Molding Department for the month of June. ROSENTHAL COMPANY Molding Department ProduCost Reconciliation Schedule Costs accounted for Transferred out Work in process, June 30 Materials Conversion costs Total co
Answer #1
a
1
Physical units
Units to be accounted for
Work in process, June 1 0
Started into production 25960
Total units 25960
Units accounted for
Transferred out 23600
Work in process, June 30 2360
Total units 25960
2
Materials Conversion Costs
Equivalent units 25960 24544
3
Materials Conversion Costs Total unit cost
Unit costs 9.00 8.00 17.00
4
Transferred out 401200
Work in process, June 30 28792
b
ROSENTHAL COMPANY
Molding Department
Production Cost Report
For the Month Ended June 30, 2020
Equivalent Units
Physical Units Materials Conversion Costs
Units to be accounted for
Work in process, June 1 0
Started into production 25960
Total units 25960
Units accounted for
Transferred out 23600 23600 23600
Work in process, June 30 2360 2360 944
Total units 25960 25960 24544
Costs Materials Conversion Costs Total
Unit costs:
Total costs 233640 196352 429992
Equivalent units 25960 24544
Unit costs 9.00 8.00 17.00
Costs to be accounted for
Work in process, June 1 0
Started into production 429992
Total costs 429992
Cost Reconciliation Schedule
Costs accounted for
Transferred out 401200 =23600*17
Work in process, June 30
      Materials 21240 =2360*9
      Conversion costs 7552 28792 =944*8
Total costs 429992